OpenStudy (anonymous):

Find the equation of the normal at x=2 to the curve with equation y=In(2x-3) I think I know how to differentiate it, but further than that, I'm stuck.

OpenStudy (anonymous):

@campbell_st can you see if you can help me, or anyone else?

OpenStudy (campbell_st):

for logs and differentiating its f'(x)/f(x) so \[dy/dx = 2/2x - 3\] substitute x = 2 to find the gradient then m = 2/(2 x 2 - 3) m = 2 substitute x = 2 into the original function to find the y value in the point y = ln(2 x 2 - 3) y = 0 since ln(1) = 0 now you have the point of contact of the tangent (2, 0) and the tangent has a gradient of 2. use the point slope equation \[y - y _{1} = m(x - _{1})\] gives y - 0 = 2( x - 2) then y = 2x - 4

OpenStudy (campbell_st):

point gradient formula \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

2y=2-x is the answer... I thought it was y=2x-4 aswell.?

OpenStudy (campbell_st):

lol... oops I thought it was the tangent.... but you need to find the normal... I should read the question more carefully. ok... for a normal the gradient is -1/m where m is the gradient of the tangent. so the normal has a gradient m= -1/2 but still passes through the point (2, 0) put them into the point gradient formula and you will get y = -x/2 + 1 multiply each term by 2 to get rid of the fraction will give 2y = 2 - x sorry about not reading the question correctly

OpenStudy (anonymous):

thanks